OpenStudy (anonymous):

Given a soda can with a volume of 15 and a diameter of 2, what is the volume of a cone that fits perfectly inside the soda can? (Hint: only enter numerals in the answer blank). Numerical Answers Expected!

OpenStudy (anonymous):

if u can give me the answer that would be great I'm sorry i normally request an explanation but i have too go too work in 10 mins

rebeccaxhawaii (rebeccaxhawaii):

got the form?

OpenStudy (anonymous):

form?

OpenStudy (anonymous):

@rebeccaxhawaii

OpenStudy (anonymous):

@haleyelizabeth2017 @kat_here2learn @geerky42 @Dani_Rose @emmagrace231 @Englishguy @Michele_Laino @purple_pink @pooja195 @Xaze @zoeyoneconovie @YanaSidlinskiy @Mr.Z @Savannah_Noelle @Adrianna.Gongora @Liv1234 @LegendaryNikki

OpenStudy (anonymous):

if you could help me thet would be amazing <3

OpenStudy (anonymous):

@Mr_Perfection_xD

OpenStudy (kat_here2learn):

5 i think

OpenStudy (kat_here2learn):

volume of cylinder=πr^2h Volume of cone=1/3πr^2h=1/3∗15=5 Because r and h are same.

OpenStudy (mr.z):

< Has most of his medals in History and English < Barely any in math and all those math questions were misplaced History and English questions. But sure, keep tagging me whatever reason.

OpenStudy (anonymous):

@Mr.Z i tagged a lot of people so if u can't answer it i don't need u being a wingspan about it just ignore it then

OpenStudy (anonymous):

wingspan*****

OpenStudy (mr.z):

Stop crying. I didn't mean to make you cry.